# Dynamics and topology of the Hat family of tilings

### v3

@@ 1. Introduction and results
This result extended earlier results of Taylor [23] and of Socolar and Taylor [20]; see also [4, Ex. 6.6] for further details. There, a functional monotile is derived from an inflation rule with decorated hexagons that defines a unique local indistinguishability (LI) class of tilings with perfect local rules [4, Def. 5.20]. Unfortunately, these local rules are represented in the form of a decoration of a hexagon that encode both nearest-neighbor and next-to-nearest-neighbor information. A purely geometric version is possible, but only with a prototile that is not connected, hence not disk-like. Both the decorated hexagon and its reflected copy are needed; they occur with equal frequency, as is nicely visible from the Llama tiling; see [4, Fig. 6.23]. The unique LI class defined by the decorated hexagon is reflection symmetric, and the tiling is limit-periodic, with pure-point spectrum, both in the dynamical and in the diffraction sense.
@@
Our first main result is that the shape changes considered in [17] are part of a $4$-dimensional family of topological conjugacies between the translation dynamical systems of the family of Hat tilings. For the connections between them, we need the concepts of local derivability and mutual local derivability (MLD); see [4, Sec. 5.2] for definitions and background.
@@
In particular, the translation dynamics of the $\mathbb{R}^{2}$ action on the hull of every version of a Hat tiling is topologically conjugate, up to linear transformation (possibly including reflection), to the translation dynamics on the hull of every other version of a Hat tiling. From a dynamical systems perspective, there is only one Hat tiling. Note that a Hat-dominated tiling is reflected into an anti-Hat dominated one, but since $\mathrm{GL}_{2}(\mathbb{R})$ is not connected, the two cannot continuously be deformed into one another.
Some of the shape changes violate the symmetries of the original construction. The family of shape changes that respect rotational symmetry is only $4$-dimensional and can be coded in two complex numbers, where we identify $\mathbb{R}^{2}$ with $\mathbb{C}$. One complex number parametrizes uniform rotations and rescalings, while the other parametrizes shape conjugacies. If we further insist that the shape changes respect reflectional symmetry, the two complex parameters must in fact be real. One represents rescalings and the other is exactly the ratio of edge lengths described in [17].
@@ 2. Shape deformations of the Hat
Given any tiling in $d$ dimensions, we can study changes to the shapes and sizes of the tiles that allow the tiles to fit together in the exact same combinatorial patterns. Some of these changes result in tilings that are MLD to the original one. Infinitesimally, shape changes modulo MLD equivalence of a tiling $\mathcal{T}$ are parametrized by $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{d})$, where $\Omega_{\mathcal{T}}$ is the orbit closure (continuous hull) of the tiling $\mathcal{T}$; see [9]. Since $d=2$ and we identify $\mathbb{R}^{2}$ with $\mathbb{C}$, we must compute the complex-valued cohomology of the hull of each tiling. There are two such hulls, depending on whether the tiling involves meta-tiles or reflected meta-tiles. Since these hulls are homeomorphic (being reflections of one another), it suffices to work with a tiling by ordinary meta-tiles.
@@
One standard procedure for computing the cohomology of a substitution tilings was developed by Anderson and Putnam [1]. We construct a branched manifold $\varGamma$, called the Anderson–Putnam (AP) complex, from one representative of each type of tile in the tiling, with appropriate edge identifications. Specifically, whenever there is a place in the tiling where tiles of type $\mathfrak{t}_{1}$ and $\mathfrak{t}_{2}$ meet along an edge, the corresponding edges of $\mathfrak{t}_{1}$ and $\mathfrak{t}_{2}$ are identified in $\varGamma$. Substitution induces a map from this complex to itself. As long as the substitution meets a technical condition called forcing the border [11] (defined below), the tiling space is homeomorphic to the inverse limit of the AP complex under substitution, and the Čech cohomology $\check{H}^{k}(\Omega_{\mathcal{T}})$ of the tiling space is the direct limit of the (ordinary) cohomology $H^{k}(\varGamma)$ of the AP complex.
Anderson and Putnam developed their machinery for self-similar tilings, but the topology of the tiling space does not depend on the geometry of the tiles. A nearly identical procedure applies to combinatorial substitution systems like the Hat, or more generally to fusion tilings of any kind; see [10, Sec. 5] for details. The $n$th instance of the AP complex is replaced by a complex constructed from $n$th order supertiles identified along supertile edges and the substitution map (between a space and itself) is replaced by a forgetful map from a complex built from $n$th order supertiles to a complex build from $(n{\hskip 0.5pt-}1)$st order supertiles.
@@
In Figure 2, we show patches of tiles which are enforced by the first order supertiles. The tiles intersecting the interior of their supertile were already given in [17, Figure 2.8]. We had to complement these patches only by a few tiles having a vertex on the supertile boundary. These extra vertices are either a purple tip of a tile or a supertile, where a local three-fold symmetry is required by the matching conditions. In the terminology of [4], we replace the inflation rule by a consistent pseudo-inflation with a complete ‘belt’ of positive thickness. These pictures show that the enforced patches extend by a positive margin beyond the perimeter of the supertile. Under inflation, the thickness of this margin, where the tiling is also determined, grows at the same rate as the supertile size, which implies border-forcing. ∎
@@
and we further need to apply the identities $L=-r^{3}L$, $X_{+}=r^{2}X_{+}$ and $F_{-}=r^{2}F_{-}$. Fortunately, it is not necessary to work with gigantic matrices. Instead, we follow the methods of [15], see also [16, Ch. 4], and work with one representation of $C_{6}$ at a time. The total cohomology is then the direct sum of the contributions of each representation.
@@ 3. The CAP tiling
This is cohomologous to $(\phi+1)\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}$, which we represent as $(\phi+1,\phi+\xi)$. This is a left-eigenvector of $\widetilde{\sigma}_{1}$ with eigenvalue $\phi^{2}$. The substitution rule (shown in Figure 2) is self-similar in the sense that a substituted tiling (with shape class $(\phi+1,\phi+\xi)\widetilde{\sigma}_{1}$) is then MLD to a rescaling of the original tiling by $\phi^{2}$ (with shape class $\phi^{2}(\phi+1,\phi+\xi)$). This constitutes the LIDS property in the sense of [4, Def. 5.16].
@@
For a stone inflation, the corresponding overlap inflation can now be used to test Solomyak’s criterion. The area fraction covered by coincident tiles converges to $1$ if and only if every overlap type eventually produces a coincidence overlap. In that case, the area fraction covered by the discrepancy overlaps shrinks at each inflation step by a certain factor bounded away from $1$. Conversely, if at least one overlap type does not produce any coincidence, there is a discrepancy region whose area fraction does not shrink.
@@
Figure 8. Window for the cut-and-project scheme of the CAP tiling. The color code refers to the type of the tile whose control point is plotted (orange for $\mathrm{T}$-, blue for $\mathrm{H}$-, green for $\mathrm{P}$- and purple for $\mathrm{F}$-tiles). The subdivision for the different tile types is partly fractal, see Eq. (17), and partly straight. The edge length of the window is $\phi$ times the unit length in $R$. The image is chiral with sixfold rotational, but no reflection symmetry, although the total window, and hence the set of all control points (without color), is mirror symmetric.
Now, we show that the set $\varLambda$ of control points of the CAP tiling, which is a subset of the return module $R$, forms a cut-and-project set (CPS). For this, we need the construction from [8] which gives the CPS from intrinsic data of the tiling; see also [22, Sec. 5.2] for a detailed description. The procedure simplifies for inflation point sets, where one can more easily determine the limit translation module from the return vectors. The upshot is that the limit translation module agrees with our return module, and needs to be lifted to an embedding lattice. Due to the underlying structure with the PV unit $\phi^{2}$, one natural way to turn $R$ into a lattice in $\mathbb{R}^{4}$ employs the classic Minkowski embedding; see [4, Sec. 3.4].
@@
Put differently, we are using the following cut-and-project scheme [4, Sec. 7.2]
@@
for $W_{i}=\overline{\varLambda_{i}^{\star}}$. Note that, due to having taken the closure, the unions on the right-hand side need no longer be disjoint. Since $\phi^{2}$ is a PV number, this IFS is contractive and defines a unique attractor $(W_{1},\ldots,W_{24})\subset(\mathcal{K}\mathbb{R}^{2})^{24}$, where $\mathcal{K}\mathbb{R}^{2}$ is the space of non-empty compact subsets of $\mathbb{R}^{2}$, equipped with the Hausdorff metric. By construction, each
@@
is a regular model set with $\varLambda_{i}\subseteq\mbox{\Large$\curlywedge$}(W_{i})$ by construction, where each window $W_{i}$ is a topologically regular compact set with almost no boundary; see [6, Sec. 3] and references therein for details. Some of the boundaries are straight lines, while some others are fractals of Koch curve type, then with Hausdorff dimension
@@
Since $\rho_{1}=\rho_{2}$, the set of control points of the inflation tiling is indeed the model set $(W)$, possibly up to a difference set of zero density. The autocorrelation measures of both sets agree, because the addition or subtraction of sets of zero density does not matter [4, Rem. 9.14], and the same applies to the diffraction of Dirac combs on $(W)$ that are weighted according to the point classes, in comparison to the correspondingly weighted Dirac combs on the control point sets. The diffraction measures of both situations are thus the same, and pure point. Since the hulls of both sets give rise to ergodic dynamical systems, the equivalence theorem [12, 7] implies that the tiling dynamical system has pure-point dynamical spectrum, and both systems have eigenfunctions with continuous representatives. In other words, topological and measure-theoretic point spectra are the same.
@@
The equality of $\rho_{1}$ and $\rho_{2}$ is also a consistency check of the correct choice of lattice for the CPS. The result now gives the group $\mathbb{R}^{4}/\mathcal{L}\simeq\mathbb{T}^{4}$ as the maximal equicontinuous factor (MEF) of our tiling space. On the level of the control points, due to the nature of the windows, the sets $\varLambda_{i}$ and $(W_{i})$ can only differ in points whose $\star$-image lies on a window boundary. If this happens (as it does here), the model set description gives the union of all elements in the fibre over the corresponding point of the MEF. They are the singular points, which have a slightly more complicated nature than usual, due to the partly fractal nature of the window boundaries. $\Diamond$
@@
The total window for all control points is a regular hexagon, whereas the window for the $\mathrm{F}$-tile control points consists of $6$ regular triangles. By the general criteria for MLD relations of regular model sets in the same CPS, see [4, Rem. 7.6], we thus see that the $\mathrm{F}$-type points can locally be derived from knowing all points (without color), and vice versa. All other subwindows partly have fractal boundaries. This means that control points of other tile types cannot be derived by local means from the (uncolored) positions of all control points. $\Diamond$
@@
which exists, for any $k\in\mathbb{R}^{2}$, due to ergodicity. Here, $B_{s}(0)$ denotes the closed disk of radius $s$ around 00. The Fourier–Bohr coefficient is an area-averaged exponential sum, with $a_{i}(0)=\mathrm{dens}(\varLambda_{i})=\mathrm{dens}(\mathcal{L})\hskip 0.5pt\mathrm{vol}(W_{i})$ from our above analysis. By the general spectral result for regular model sets, compare [4, Thm. 9.4], we get
@@
While the formula in (19) is a major result in the theory of regular model sets, its calculation can become complicated, in particular for windows with fractal boundaries. Here, the windows $W_{i}$ are still such that a numerical approach (as in [19]) gives reasonable approximations. An exact calculation is also possible, via the compactly converging Fourier cocycle developed in [6, Sec. 4], which works here in exactly the same way. A simple example, also in comparison to a numerical approximation, is worked out in [5]. $\Diamond$
@@ 3.1. Reprojections
is another projection from $\mathbb{R}^{4}$ to $\mathbb{R}^{2}$. In fact, every projection from the total space $\mathbb{R}^{4}$ to direct space $\mathbb{R}^{2}$ can be written in this way. Changing projections is exactly the same as adding a linear map from internal space to direct space, which is a standard situation in the treatment of deformed model sets; see [4, Ex. 9.9] and references given there.
@@
By reprojecting the CAP Delone set in arbitrary ways, we obtain Delone sets whose shape class differ from the CAP shape class by arbitrary asymptotically negligible elements of $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{2})$. This means that every tiling that is topologically conjugate to the CAP Delone set is MLD to such a reprojection of the CAP Delone set, as claimed in Theorem 3.
@@
Under topological conjugacy, singular points of the MEF must be mapped onto each other, preserving the cardinality of the fibres. The set of points in the MEF having a singular fibre (with more than one element) is given by $S=(\mathbb{R}^{2}\times\partial W)/\mathcal{L}$, where $\partial W$ is the set of all boundary points of the window (including internal boundaries), and $\mathbb{R}^{2}$ are the translations in the direction of direct space. This is a subset of measure zero of the MEF $(\mathbb{R}^{2}\times\mathbb{R}^{2})/\mathcal{L}$. Due to the fractal boundaries, a reflection preserving the total window cannot also preserve the set $S$. This means that translations of the hull in direct space direction cannot commute with a reflection, for any direction of projection, even if they do so for the MEF itself. Hence, a reflection can preserve the spectrum under certain conditions, but it cannot be a topological conjugacy. $\Diamond$

### v4

@@ 1. Introduction and results
This result extended earlier results of Taylor [24] and of Socolar and Taylor [21]; see also [5, Ex. 6.6] for further details. There, a functional monotile is derived from an inflation rule with decorated hexagons that defines a unique local indistinguishability (LI) class of tilings with perfect local rules [5, Def. 5.20]. Unfortunately, these local rules are represented in the form of a decoration of a hexagon that encodes both nearest-neighbor and next-to-nearest-neighbor information. A purely geometric version is possible, but only with a connected prototile that is not simply connected, hence not disk-like. Both the decorated hexagon and its reflected copy are needed; they occur with equal frequency, as is nicely visible from the Llama tiling; see [5, Fig. 6.23]. The unique LI class defined by the decorated hexagon is reflection symmetric, and the tiling is limit-periodic, with pure-point spectrum, both in the dynamical and in the diffraction sense.
@@
Our first main result is that the shape changes considered in [18] are part of a $4$-dimensional family of topological conjugacies between the translation dynamical systems of the family of Hat tilings. For the connections between them, we need the concepts of local derivability and mutual local derivability (MLD); see [5, Sec. 5.2] for definitions and background.
@@
In particular, the translation dynamics of the $\mathbb{R}^{2}$ action on the hull of every version of a Hat tiling is topologically conjugate, up to linear transformation (possibly including reflection), to the translation dynamics on the hull of every other version of a Hat tiling. From a dynamical systems perspective, there is only one Hat tiling. Note that a Hat-dominated tiling is reflected into an anti-Hat dominated one, but since $\mathrm{GL}_{2}(\mathbb{R})$ is not connected, the two cannot be deformed continuously into one another.
Some of the shape changes violate the symmetries of the original construction. The family of shape changes that respect rotational symmetry is only $4$-dimensional and can be coded in two complex numbers, where we identify $\mathbb{R}^{2}$ with $\mathbb{C}$. One complex number parameterizes uniform rotations and rescalings, while the other parameterizes shape conjugacies. If we further insist that the shape changes respect reflectional symmetry, the two complex parameters must in fact be real. One represents rescalings and the other is exactly the ratio of edge lengths described in [18].
@@ 2. Shape deformations of the Hat
Given any tiling in $d$ dimensions, we can study changes to the shapes and sizes of the tiles that allow the tiles to fit together in the exact same combinatorial patterns. Some of these changes result in tilings that are MLD to the original one. Infinitesimally, shape changes modulo MLD equivalence of a tiling $\mathcal{T}$ are parameterized by $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{d})$, where $\Omega_{\mathcal{T}}$ is the orbit closure (continuous hull) of the tiling $\mathcal{T}$; see [10]. Since $d=2$ and we identify $\mathbb{R}^{2}$ with $\mathbb{C}$, we must compute the complex-valued cohomology of the hull of each tiling. There are two such hulls, depending on whether the tiling involves meta-tiles or reflected meta-tiles. Since these hulls are homeomorphic (being reflections of one another), it suffices to work with a tiling by ordinary meta-tiles.
@@
One standard procedure for computing the cohomology of a substitution tiling was developed by Anderson and Putnam [1]. We construct a branched manifold $\varGamma$, called the Anderson–Putnam (AP) complex, from one representative of each type of tile in the tiling, with appropriate edge identifications. Specifically, whenever there is a place in the tiling where tiles of type $\mathfrak{t}_{1}$ and $\mathfrak{t}_{2}$ meet along an edge, the corresponding edges of $\mathfrak{t}_{1}$ and $\mathfrak{t}_{2}$ are identified in $\varGamma$. Substitution induces a map from this complex to itself. As long as the substitution meets a technical condition called forcing the border [12] (defined below), the tiling space is homeomorphic to the inverse limit of the AP complex under substitution, and the Čech cohomology $\check{H}^{k}(\Omega_{\mathcal{T}})$ of the tiling space is the direct limit of the (ordinary) cohomology $H^{k}(\varGamma)$ of the AP complex.
Anderson and Putnam developed their machinery for self-similar tilings, but the topology of the tiling space does not depend on the geometry of the tiles. A nearly identical procedure applies to combinatorial substitution systems like the Hat, or more generally to fusion tilings of any kind; see [11, Sec. 5] for details. The $n$th instance of the AP complex is replaced by a complex constructed from $n$th order supertiles identified along supertile edges and the substitution map (between a space and itself) is replaced by a forgetful map from a complex built from $n$th order supertiles to a complex build from $(n{\hskip 0.5pt-}1)$st order supertiles.
@@
In Figure 2, we show patches of tiles which are enforced by the first order supertiles. The tiles intersecting the interior of their supertile were already given in [18, Figure 2.8]. We had to complement these patches only by a few tiles having a vertex on the supertile boundary. These extra vertices are either a purple tip of a tile or a supertile, where a local three-fold symmetry is required by the matching conditions. In the terminology of [5], we replace the inflation rule by a consistent pseudo-inflation with a complete ‘belt’ of positive thickness. These pictures show that the enforced patches extend by a positive margin beyond the perimeter of the supertile. Under inflation, the thickness of this margin, where the tiling is also determined, grows at the same rate as the supertile size, which implies border-forcing. ∎
@@
and we further need to apply the identities $L=-r^{3}L$, $X_{+}=r^{2}X_{+}$ and $F_{-}=r^{2}F_{-}$. Fortunately, it is not necessary to work with gigantic matrices. Instead, we follow the methods of [16], see also [17, Ch. 4], and work with one representation of $C_{6}$ at a time. The total cohomology is then the direct sum of the contributions of each representation.
@@ 3. The CAP tiling
This is cohomologous to $(\phi+1)\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}$, which we represent as $(\phi+1,\phi+\xi)$. This is a left-eigenvector of $\widetilde{\sigma}_{1}$ with eigenvalue $\phi^{2}$. The substitution rule (shown in Figure 2) is self-similar in the sense that a substituted tiling (with shape class $(\phi+1,\phi+\xi)\widetilde{\sigma}_{1}$) is then MLD to a rescaling of the original tiling by $\phi^{2}$ (with shape class $\phi^{2}(\phi+1,\phi+\xi)$). This constitutes the LIDS property in the sense of [5, Def. 5.16].
@@
For a stone inflation, the corresponding overlap inflation can now be used to test Solomyak’s criterion. The area fraction covered by coincident tiles converges to $1$ if and only if every overlap type eventually produces a coincidence overlap. In this case, the area fraction covered by the discrepancy overlaps shrinks at each inflation step by a certain factor bounded away from $1$. Conversely, if at least one overlap type does not produce any coincidence, there is a discrepancy region whose area fraction does not shrink.
@@
Figure 8. Window for the cut-and-project scheme of the CAP tiling. The color code refers to the type of the tile whose control point is plotted (orange for $\mathrm{T}$-, blue for $\mathrm{H}$-, green for $\mathrm{P}$- and purple for $\mathrm{F}$-tiles). The subdivision for the different tile types is partly fractal, see Eq. (17), and partly straight. The edge length of the window is $\phi$ times the unit length in $R$. The image is chiral with sixfold rotational symmetry, but has no reflection symmetry, although the total window, and hence the set of all control points (without color), is mirror symmetric.
Now, we show that the set $\varLambda$ of control points of the CAP tiling, which is a subset of the return module $R$, forms a cut-and-project set (CPS). For this, we need the construction from [9] which gives the CPS from intrinsic data of the tiling; see also [23, Sec. 5.2] for a detailed description. The procedure simplifies for inflation point sets, where one can more easily determine the limit translation module from the return vectors. The upshot is that the limit translation module agrees with our return module, and needs to be lifted to an embedding lattice. Due to the underlying structure with the PV unit $\phi^{2}$, one natural way to turn $R$ into a lattice in $\mathbb{R}^{4}$ employs the classic Minkowski embedding; see [5, Sec. 3.4].
@@
Put differently, we are using the following cut-and-project scheme [5, Sec. 7.2]
@@
for $W_{i}=\overline{\varLambda_{i}^{\star}}$. Note that, due to having taken the closure, the unions on the right-hand side need no longer be disjoint. Since $\phi^{2}$ is a PV number, this IFS is contractive and defines a unique attractor $(W_{1},\ldots,W_{24})\subset(\mathcal{K}\mathbb{R}^{2})^{24}$, where $\mathcal{K}\mathbb{R}^{2}$ is the space of non-empty compact subsets of $\mathbb{R}^{2}$, equipped with the Hausdorff metric. Each
@@
is a regular model set with $\varLambda_{i}\subseteq\mbox{\Large$\curlywedge$}(W_{i})$ by construction, where each window $W_{i}$ is a topologically regular compact set with almost no boundary; see [7, Sec. 3] and references therein for details. Some of the boundaries are straight lines, while some others are fractals of Koch curve type, then with Hausdorff dimension
@@
Since $\rho_{1}=\rho_{2}$, the set of control points of the inflation tiling is indeed the model set $(W)$, possibly up to a difference set of zero density. The autocorrelation measures of both sets agree, because the addition or subtraction of sets of zero density does not matter [5, Rem. 9.14], and the same applies to the diffraction of Dirac combs on $(W)$ that are weighted according to the point classes, in comparison to the correspondingly weighted Dirac combs on the control point sets. The diffraction measures of both situations are thus the same, and pure point. Since the hulls of both sets give rise to ergodic dynamical systems, the equivalence theorem [13, 8] implies that the tiling dynamical system has pure-point dynamical spectrum, and both systems have eigenfunctions with continuous representatives. In other words, topological and measure-theoretic point spectra are the same.
@@
The equality of $\rho_{1}$ and $\rho_{2}$ is also a consistency check of the correct choice of lattice for the CPS. The result now gives the group $\mathbb{R}^{4}/\mathcal{L}\simeq\mathbb{T}^{4}$ as the maximal equicontinuous factor (MEF) of our tiling space. On the level of the control points, due to the nature of the windows, the sets $\varLambda_{i}$ and $(W_{i})$ can only differ in points whose $\star$-image lies on a window boundary. If this happens (as it does here), the model set description gives the union of all elements in the fiber over the corresponding point of the MEF. They are the singular points, which have a slightly more complicated nature than usual, due to the partly fractal nature of the window boundaries. $\Diamond$
@@
The total window for all control points is a regular hexagon, whereas the window for the $\mathrm{F}$-tile control points consists of $6$ regular triangles. By the general criteria for MLD relations of regular model sets in the same CPS, see [5, Rem. 7.6], we thus see that the $\mathrm{F}$-type points can locally be derived from knowing all points (without color), and vice versa. All other subwindows partly have fractal boundaries. This means that control points of other tile types cannot be derived by local means from the (uncolored) positions of all control points. $\Diamond$
@@
which exists, for any $k\in\mathbb{R}^{2}$, due to ergodicity. Here, $B_{s}(0)$ denotes the closed disk of radius $s$ around 00. The Fourier–Bohr coefficient is an area-averaged exponential sum, with $a_{i}(0)=\mathrm{dens}(\varLambda_{i})=\mathrm{dens}(\mathcal{L})\hskip 0.5pt\mathrm{vol}(W_{i})$ from our above analysis. By the general spectral result for regular model sets, compare [5, Thm. 9.4], we get
@@
While the formula in (19) is a major result in the theory of regular model sets, its calculation can become complicated, in particular for windows with fractal boundaries. Here, the windows $W_{i}$ are still such that a numerical approach (as in [20]) gives reasonable approximations. An exact calculation is also possible, via the compactly converging Fourier cocycle developed in [7, Sec. 4], which works here in exactly the same way. A simple example, also in comparison to a numerical approximation, is worked out in [6]. $\Diamond$
@@ 3.1. Reprojections
is another projection from $\mathbb{R}^{4}$ to $\mathbb{R}^{2}$. In fact, every projection from the total space $\mathbb{R}^{4}$ to direct space $\mathbb{R}^{2}$ can be written in this way. Changing projections is exactly the same as adding a linear map from internal space to direct space, which is a standard situation in the treatment of deformed model sets; see [5, Ex. 9.9] and references given there.
@@
By reprojecting the CAP Delone set in arbitrary ways, we obtain Delone sets whose shape classes differ from the CAP shape class by arbitrary asymptotically negligible elements of $\check{H}^{1}(\Omega_{\mathcal{T}},\mathbb{R}^{2})$. This means that every tiling that is topologically conjugate to the CAP Delone set is MLD to such a reprojection of the CAP Delone set, as claimed in Theorem 3.
@@
Under topological conjugacy, singular points of the MEF must be mapped onto each other, preserving the cardinality of the fibers. The set of points in the MEF having a singular fiber (with more than one element) is given by $S=(\mathbb{R}^{2}\times\partial W)/\mathcal{L}$, where $\partial W$ is the set of all boundary points of the window (including internal boundaries), and $\mathbb{R}^{2}$ are the translations in the direction of direct space. This is a subset of measure zero of the MEF $(\mathbb{R}^{2}\times\mathbb{R}^{2})/\mathcal{L}$. Due to the fractal boundaries, a reflection preserving the total window cannot also preserve the set $S$. This means that translations of the hull in direct space direction cannot commute with a reflection, for any direction of projection, even if they do so for the MEF itself. Hence, a reflection can preserve the spectrum under certain conditions, but it cannot be a topological conjugacy. $\Diamond$